Find the net rate of radiation by a pot at 46.1 ∘C that has been placed in a −20.7 ∘C freezer. The pot's surface area is 0.131 m2and its emissivity is 0.581.

net rate of radiation:
Answer #1

Stefan-Boltzmann constant = = 5.67 x 10-8 W/(m2.K4)

Surface area of the pot = A = 0.131 m2

Emissivity of the pot = e = 0.581

Temperature of the pot = T1 = 46.1 oC = 46.1 + 273 K = 319.1 K

Temperature of the freezer = T2 = -20.7 oC = -20.7 + 273 K = 252.3 K

Net rate of radiation from the pot = Q

Q = Ae(T14 - T24)

Q = (0.131)(0.581)(5.67x10-8)(319.14 - 252.34)

Q = 27.26 W

Net rate of radiation from the pot = 27.26 W